# Dissecting CLIP: Decomposition with a Schur Complement-based Approach

### v1

@@ 1 Introduction
Figure 2: Clusters of DALL-E 3 [1] generated images of animals with fruit-labeled tags. The yellow column shows Kernel-PCA (KPCA) clusters using CLIP; the green column shows KPCA clusters with the application the proposed Schur-Complement-based method to remove "fruit" direction from CLIP embedding; and the blue column shows clusters with the "animal" direction removed from CLIP embedding.
@@
To achieve this goal, we employ a kernel similarity function and focus on the kernel covariance matrix for the generated (text, image) pairs. Using a Gaussian elimination-based approach in a proper kernel space, we decompose the image sub-covariance matrix, $C_{II}$, into the sum of the Schur complement component, $C_{II}-C_{IT}C_{TT}^{-1}C_{TI}$, which represents the model-induced diversity component, and the remainder term $C_{IT}C_{TT}^{-1}C_{TI}$, which captures the text-induced variety in the generated images. Leveraging the estimated kernel covariance matrices from generated data, we define the matrix-based entropy of these components as measures of model-induced and text-induced diversity in the generated images. Specifically, we define the Schur Complement Entropy (SCE) score as the matrix-based entropy of the non-text-based component to measure the intrinsic diversity of images.
Additionally, we use the eigenspace of the Schur complement component to interpret the diversity in the generated images that is independent of the text prompt, revealing the unique elements introduced by the generation model in the images. The eigenspace-based interpretation follows the application of the Kernel PCA (KPCA) clustering method and leads to a visualization of the sample clusters due to the input text clusters and the clusters with a feature added by the model to the generated image. Figure 2 shows the resulting KPCA-detected clusters for DALL-E 3 [1] generated images in response to prompts of type "[animal] eating a [fruit]" with 5 different animal and fruit names. The detected clusters of CLIP image embedding (leftmost case) shows clusters with mixed animals and fruits. On the other hand, by considering the prompt of only specifying the fruit, the Schur complement component’s KPCA clusters highlight the animals in each cluster (middle case), and similarly by considering the prompt of only specifying the fruit, the animals are captured by each KPCA cluster of the Schur complement matrix.
@@
We present several numerical applications of the proposed framework to evaluate and interpret standard text-to-image and image-captioning models. Our results on several simulated scenarios with known ground-truth diversity indicate that the proposed entropy metric correlates with the non-text-based diversity in images, capturing variation not attributable to the text prompt. Additionally, we show that the decomposed feature component can neutralize the influence of specific objects in the text prompt within the image embedding. Specifically, we use this decomposed feature to diminish the impact of visible text in images, reducing its effect on the embedding. We also demonstrate how the decomposed image embedding can enhance or reduce the focus on particular objects or styles within an image, which can have implications for downstream applications of CLIP embeddings when emphasizing specific elements. For instance, Figure 1 shows a generated image of two random concepts. We demonstrate the ImageNet samples with highest CLIPScore to the original image embedding and to the modified image embedding when subtracting the Schur complement-method embedding after canceling corresponding appearing concepts as text, such as "pumpkin" and "ostrich", which shows the modified embedding removes the effect of the unwanted object. The main contributions of this work are summarized as follows:
@@ 3 Preliminaries
In this work, we provide several numerical results results for the cosine-similarity Kernel function defined as:
@@ 4 Diversity Evaluation for Text-to-Image Generative Models via CLIP Embedding
Proposition 1 shows that given the optimal text-to-image conversion matrix $\Gamma^{*}=C_{IT}C_{TT}^{-1}$, the effect of a text $T$ on the embedding of an image $I$ can be canceled by considering the remainder term $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$ to decorrelate the image and text embedding.
Following the above discussion for the cosine-similarity kernel $\phi(x)=x/\|x\|$, we can modify the CLIP embedding given a paired dataset of prompts and images. To do this, we first compute the modification matrix $\Gamma^{*}=C_{IT}C_{TT}^{-1}$ and modify the CLIP embedding for an image sample $x_{I}=\mathrm{CLIP}(I)$ and correcting prompt $x_{T}=\mathrm{CLIP}(T)$ as:
@@
(a) Unconditional Vendi and conditional SCE entropy scores for SDXL-generated cats from different breeds. We used two sets of prompts for diversity evaluation: (upper) unspecified cat breed (lower) specified cat breed
(b) Unconditional Vendi and conditional SCE entropy scores for SDXL-generated animals with objects. We used two sets of prompts for diversity evaluation: (upper) specified animals only (lower) specified animals and objects
Figure 6: Evaluated SCE and Vendi scores with Gaussian Kernel on two datasets: different cat breeds and animals with objects.
Following the above decomposition, we define the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$ and the text part $\mathrm{SCE_{T}}$ as follows:
where $\lambda^{(\Lambda_{I})}_{j}$ denotes the $j$th eigenvalue of matrix $\Lambda_{I}$ and $\mathrm{Tr}(\Lambda_{I})=\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}$ is the sum of the eigenvalues. Note that we follow the same definition for the text part $\Lambda_{T}$. It can be seen that the defined Schur complement entropy is the conditional entropy of the hidden image cluster variable $\mathrm{Mode}\in\{1,2,\ldots,d\}$, which follows the spectral decomposition of the image covariance matrix $C_{II}=\sum_{j=1}^{d}\lambda_{j}v_{j}v_{j}^{\top}$, where the mode $\mathrm{Mode}=i$ has probability $\lambda_{i}$. Here we define a text-based guess of the image mode $Y_{T}$, which as we discuss in the Appendix can correctly predict the cluster with probability $\mathrm{Tr}(\Lambda_{T})$ and outputs erasure $e$ with probability $\mathrm{Tr}(\Lambda_{I})$. Then, we show in the Appendix that following standard Shannon entropy definition:
$\mathrm{SCE}_{I}(x_{1},\ldots,x_{n})=H\bigl{(}\mathrm{Mode}|Y_{T}\bigr{)}$
Therefore, this analysis allows us to decompose the diversity of generated images into model-induced and text-induced components.
5 Numerical Results
Figure 7: t-SNE plot of animals with objects dataset.
@@ 5 Numerical Results
We evaluated our proposed Schur Complement-based approach for CLIP decomposition across various scenarios using both real and synthetic datasets. Our experimental results are reported for the standard Cosine Similarity and Gaussian kernel functions. For the Gaussian kernel function, we use random Fourier features [26] with a random Fourier dimension $r=2000$ to embed the kernel using a finite feature dimension. In the diversity evaluation, we report our defined SCE scores as well as the Vendi score [7] for unconditional diversity assessment without taking the text data into account. The clustering experiments were performed using Kernel PCA with the Gaussian kernel, highlighting the top clusters that align with the top eigenvector directions in the data. For CLIP correction experiments, we computed correction matrix $\Gamma^{*}$ on MSCOCO [27] train dataset and used the matrix in our experiments.
Measuring Text-to-Image Model Diversity. To evaluate text-to-image model diversity, we decompose the unconditional CLIP-based Vendi score [7], i.e., the unconditional entropy of images, into two components: $\exp(\mathrm{SCE}_{I})$ score for the model-induced diversity in images and $\exp(\mathrm{SCE}_{T})$ score for the text-induced diversity. Figure 6 illustrates this diversity decomposition on a dataset of generated images using Gaussian Kernel on top of the CLIP embedding. In this setup, we generated 1,000 images per class using SDXL [28] and plotted the decomposed metrics. When prompts do not specify cat breeds, text-based diversity $\mathrm{SCE}_{T}$ remains almost constant, and image diversity $\mathrm{SCE}_{I}$ is close to the unconditional Vendi score. Conversely, when prompts specify cat breeds, the diversity mainly stems from the text prompts rather than the image generator, resulting in a constant $\mathrm{SCE}_{I}$ score. This observation holds true when we switch to dataset with higher number of subjects, rather than features. We constructed a dataset that contains one object from {‘boat’, ‘skyscraper’, ‘basketball’, ‘couch’, ‘umbrella’, ‘sunflowers’} paired with one animal from {‘cat’, ‘dog’, ‘guinea pig’, ‘snake’}. When prompts only specify animals, $\mathrm{SCE}_{I}$ follows the Vendi score; however, when we specify animals and objects, the diversity of image generator remains constant. The score convergence holds true across different kernel functions. Additional experiments on other cases are discussed in the Appendix.
@@
To further support our findings, we performed dimensionality reduction for CLIP and corrected CLIP embeddings using the t-SNE method. Distinct clusters of objects and animals merge when we remove either "animal" or "object" directions from the embedding, illustrating the effect of our decomposition. Figure 7 illustrates the change in embedding space after the Schur-complement-based decomposition.
@@
Typographic Attacks. CLIP’s sensitivity to text within images, as showed in [13], makes it vulnerable to typographic attacks, where overlaying misleading text on an image influences its classification [31]. To investigate this, we constructed a dataset of 10 ImageNet classes, each image engraved with a random class label. Figure 8 demonstrates CLIP’s susceptibility, showing that top eigenvector directions capture the engraved text rather than the actual image content. By conditioning on prompts like "text reading ’cassette player’ on top of image" and removing this direction, we re-clustered images based on the corrected CLIP embedding. This adjustment allowed us to identify clusters based on the actual image content rather than the engraved text. Additional results are provided in the Appendix.

### v2

@@ Abstract
Figure 2: Comparison of model-driven (left) and prompt-driven (right) diversity in 200 images generated for similar "cat and dog" prompts (left) vs. 200 images generated for diverse "cat and dog" prompts (right). While unconditional metrics favor the right-side case, the SCE score isolates intrinsic model diversity, suggesting greater model diversity on the left.
@@ 1 Introduction
To achieve this goal, we employ a kernel similarity function and focus on the kernel covariance matrix for the generated (text, image) pairs. Using a Gaussian elimination-based approach in a proper kernel space, we decompose the image sub-covariance matrix, $C_{II}$, into the sum of the Schur complement component, $C_{II}-C_{IT}C_{TT}^{-1}C_{TI}$, which represents the model-driven diversity component, and the remainder term $C_{IT}C_{TT}^{-1}C_{TI}$, which captures the prompt-driven variety in the generated images. Leveraging the estimated kernel covariance matrices from generated data, we define the matrix-based entropy of the Schur complement component, as Schur Complement Entropy (SCE), as a measure of model-driven diversity in the generated images.
To illustrate the effectiveness of the SCE score in distinguishing model-driven diversity from prompt-driven effects, we present a comparative example. Figure 2 compares the SCE-Image model diversity scores for samples generated by DALL-E 3 [7]. The left-side subfigure represents model-driven diversity, showing images of diverse-breed cats and dogs generated using similar prompts like “cat and dog read a text.” The right-side subfigure illustrates prompt-driven diversity, featuring same-breed cats and dogs across diverse-occasion prompts. While unconditional diversity metrics such as Vendi and RKE report higher diversity for the right-side case, the SCE score, designed to isolate intrinsic model diversity from prompt-driven effects, indicates greater diversity in the left-side case. This example highlights the advantage of the prompt-conditioned SCE score over existing unconditional diversity metrics.
@@
We present several numerical applications of the proposed framework to evaluate and interpret standard text-to-image and image-captioning models. Our results on several simulated scenarios with known ground-truth diversity indicate that the proposed entropy metric correlates with the non-text-based diversity in images, capturing variation not attributable to the text prompt. Additionally, we show that the decomposed feature component can neutralize the influence of specific objects in the text prompt within the image embedding. Specifically, we use this decomposed feature to diminish the impact of visible text in images, reducing its effect on the embedding. We also demonstrate how the decomposed image embedding can enhance or reduce the focus on particular objects or styles within an image, which can have implications for downstream applications of CLIP embeddings when emphasizing specific elements. For instance, Figure 1 shows a generated image of an ostrich next to a pumpkin. We demonstrate the ImageNet samples with highest CLIPScore to the original image embedding and to the modified image embedding when subtracting the Schur complement-method embedding after canceling texts "pumpkin" and "ostrich", which shows the modified embedding removes the effect of the unwanted object. The main contributions of this work are summarized as follows:
@@ 3 Preliminaries
In this work, we provide several numerical results for the cosine-similarity Kernel function defined as:
@@ 4 Diversity Evaluation for Text-to-Image Generative Models via CLIP Embedding
Proposition 1 shows that given the optimal text-to-image conversion matrix $\Gamma^{*}=C_{TI}C_{TT}^{-1}$, the effect of a text $T$ on the embedding of an image $I$ can be canceled by considering the remainder term $\phi(x_{I})-\Gamma^{*}\phi(x_{T})$ to decorrelate the image and text embedding.
Following the above discussion for the cosine-similarity kernel $\phi(x)=x/\|x\|$, we can modify the CLIP embedding given a paired dataset of prompts and images. To do this, we first compute the modification matrix $\Gamma^{*}=C_{TI}C_{TT}^{-1}$ and modify the CLIP embedding for an image sample $x_{I}=\mathrm{CLIP}(I)$ and correcting prompt $x_{T}=\mathrm{CLIP}(T)$ as:
@@
Following the above decomposition, we define the Schur-Complement-Entropy for the image part $\mathrm{SCE_{I}}$ as follows:
where $\lambda^{(\Lambda_{I})}_{j}$ denotes the $j$th eigenvalue of matrix $\Lambda_{I}$ and $\mathrm{Tr}(\Lambda_{I})=\sum_{j=1}^{d}\lambda^{(\Lambda_{I})}_{j}$ is the sum of the eigenvalues. It can be seen that the defined Schur complement entropy is the conditional entropy of the hidden image cluster variable $\mathrm{Mode}\in\{1,2,\ldots,d\}$, which follows the spectral decomposition of the image covariance matrix $C_{II}=\sum_{j=1}^{d}\lambda_{j}v_{j}v_{j}^{\top}$, where the mode $\mathrm{Mode}=i$ has probability $\lambda_{i}$. Here we define a text-based guess of the image mode $Y_{T}$, which as we discuss in the Appendix can correctly predict the cluster with probability $\mathrm{Tr}(\Lambda_{T})$ and outputs erasure $e$ with probability $\mathrm{Tr}(\Lambda_{I})$. Then, we show in the Appendix that following standard Shannon entropy definition:
$\mathrm{SCE}_{I}(x_{1},\ldots,x_{n})=H\bigl{(}\mathrm{Mode}|Y_{T}\bigr{)}$
Therefore, this analysis allows us to decompose the diversity of generated images into model-driven and text-driven components.
5 Numerical Results
Figure 3: Evaluated $\mathrm{SCE_{I}}$ scores with fixed text prompt distribution about doctors in various settings. Interpolated surface plot in the bottom right visualises the change in diversity (z-axis) when doctors have a varying diversity of age and gender features. (x and y axis)
We evaluated our proposed Schur Complement-based approach for CLIP decomposition across various scenarios using both real and synthetic datasets. Our experimental results are reported for the standard Cosine Similarity and Gaussian kernel functions. For the Gaussian kernel function, we use random Fourier features [26] with a random Fourier dimension $r=2000$ to embed the kernel using a finite feature dimension. In the diversity evaluation, we report our defined SCE scores as well as the Vendi score [6], RKE[18], Recall[4] and Coverage[5] for unconditional diversity assessment without taking the text data into account. Note that Recall and Coverage may only be performed in the presence of a valid representative reference dataset. The clustering experiments were performed using Kernel PCA with the Gaussian kernel, highlighting the top clusters that align with the top eigenvector directions in the data.
Measuring Text-to-Image Model Diversity. To highlight the strengths of SCE-based diversity evaluation, we conducted numerical experiments comparing unconditional and conditional diversity assessments. Figure 3 presents an experiment with the FLUX.1-dev [27] model, where individuals are generated based on an ambiguous prompt. We fixed the text prompt distribution and conditioned the model to generate doctors of varying gender and age. The results show that when the model is restricted to producing only "senior male doctors," the $\mathrm{SCE_{I}}$ score is the lowest. Conversely, when the model generates a broader range of age groups and genders, the diversity score increases. We also provide an interpolated surface plot illustrating the intermediate mixing results. A diversity level of 0 indicates a strong bias toward a single characteristic (e.g., all images depict males), whereas a score of 1 signifies an equal representation of all characteristics (e.g., an equal number of male and female images).
Figure 4: Evaluated $\mathrm{SCE_{I}}$ scores with fixed text prompt distribution. $\mathrm{SCE_{I}}$ changes to reflect the diversity contributed by the varying level of information in text prompts.
Figure 4 illustrates the inverse scenario: the image distribution is fixed while the accompanying text prompts vary in their specificity. We generated four sets of "junior male doctor" images with different seeds, while prompts were assigned four levels of detail—fully specified (age and gender), partially specified (age only or gender only), or unspecified (no details). We measured $\mathrm{SCE_{I}}$, Vendi, RKE, Coverage and Recall scores across four scenarios. Reference set for R/C is "doctor" class in IdenProf dataset [28]. Vendi, RKE, Recall and Coverage scores remain relatively unchanged, whereas $\mathrm{SCE_{I}}$ distinguishes between these cases. When the description is incomplete, $\mathrm{SCE_{I}}$ is lowest due to the strong presence of "junior male doctors" across all images, indicating bias. However, when prompts explicitly state that all images should depict "junior male doctors," diversity increases and is evaluated based on auxiliary features such as background details rather than age or gender. When partial information is given, $\mathrm{SCE_{I}}$ falls between the fully specified and unspecified cases. These results highlight the need for differentiation between model-driven and prompt-driven diversity.
To further quantify our findings, in Figure 5 we conducted an experiment on SDXL generated cat images of different breeds. We generated 1,000 images per class and compared $\mathrm{SCE_{I}}$ with other diversity metrics. When prompts did not specify cat breeds, the diversity captured by $\mathrm{SCE_{I}}$ aligns with unconditional Vendi score. Conversely, when breed information was included in the prompts, diversity mainly stemmed from the text prompts rather than the image generator, leading to a non-increasing $\mathrm{SCE_{I}}$ score. In both scenarios, Vendi, Coverage, and Recall scores remained unchanged. The reference set for Recall and Coverage consisted of ImageNet [29] ’cat’ samples. Additional experiments on other cases are discussed in the Appendix.
Figure 5: Evaluated $\mathrm{SCE_{I}}$, Vendi, Recall and Coverage scores with Gaussian Kernel on different cat breeds dataset.
Figure 6: t-SNE plot of animals with objects dataset.
The presented experimental results highlight the importance of disentangling the sources of diversity in text-to-image generation—specifically, the diversity induced by the text prompts versus that originating from the generative model itself. Our findings reveal that when the distribution of prompts or the resulting images is varied, a reliable evaluation metric must be sensitive to both aspects, adapting accordingly to reflect shifts in either distribution. In contrast, existing metrics primarily assess diversity based solely on image features, overlooking the influence of prompt variability. This limitation reduces their effectiveness in scenarios where prompt semantics play a central role, such as in text-to-image generation. These observations motivate the development of a new metric that jointly captures the interplay between prompt diversity and image diversity, enabling more faithful and nuanced evaluation of generative models in this setting.
@@ 5 Numerical Results
Figure 8: Clusters of DALL-E 3 [7] generated images of animals with fruits. The yellow column shows Kernel-PCA (KPCA) clusters using CLIP; the green column shows KPCA clusters with the application the proposed Schur-Complement-based method to remove "fruit" direction from CLIP embedding; and the blue column shows clusters with the "animal" direction removed from CLIP embedding.
Figure 10: Effect of removing encoded "cassette player" text on top of ImageNet samples.
@@
CLIP Decomposition. To illustrate the effect of Schur-complement-based decomposition on the embedding space, we first generate an "animals with objects" dataset with SDXL, where each sample contains one object from ‘boat’, ‘skyscraper’, ‘basketball’, ‘couch’, ‘umbrella’, ‘sunflowers’ paired with one animal from ‘cat’, ‘dog’, ‘guinea pig’, ‘snake’. In this dataset, diversity increases primarily due to the higher number of subjects rather than feature variations. Using this dataset, we apply t-SNE to visualize the dimensionality-reduced embeddings of CLIP and its corrected variant in Figure 6. The results show that distinct clusters of objects and animals merge when we remove either the "animal" or "object" direction from the embedding, demonstrating the impact of our decomposition. Performing $\mathrm{SCE}$ decomposition causes a change in the embedding space and correlation of features along removed direction greatly decreases.
@@
Figure 7b further extends our analysis by applying the $\mathrm{SCE}$-based decomposition not only to the text-to-image setting but also to its inverse: the image-to-text, or image captioning, task. The experimental setup mirrors that of Figure 7a, with the key difference being that, instead of generating images for text prompts, we generated captions for the corresponding images. For this task, we used gpt4o-mini as the captioning model. We selected images closely aligned with the concept we aimed to remove. For instance, to eliminate the "cat" direction in text, we used an image of a cat against a white background to better isolate the concept. After applying corrections for "animals" or "objects" in the text prompt, we observed successful decomposition, as reflected in the second column: the corrected CLIP embedding is no longer sensitive to the removed concept. The results closely mirror those observed in the text-to-image case, reinforcing that our decomposition method consistently removes the influence of the specified concept across both textual and visual modalities. This symmetry highlights the robustness of our approach in isolating concept-driven variations, regardless of the generation direction. Our method can be readily extended to other paired tasks, provided that an appropriate embedding model exists to encode the paired modalities.
Additionally, we use the eigenspace of the Schur complement component to interpret the diversity in the generated images that is independent of the text prompt, revealing the unique elements introduced by the generation model in the images. The eigenspace-based interpretation follows the application of the Kernel PCA (KPCA) clustering method and leads to a visualization of the sample clusters due to the input text clusters and the clusters with a feature added by the model to the generated image. Figure 8 shows the resulting KPCA-detected clusters for DALL-E 3 [7] generated images in response to prompts of type "[animal] eating a [fruit]" with 5 different animal and fruit names. The detected clusters of CLIP image embedding (leftmost case) shows clusters with mixed animals and fruits. On the other hand, by considering the prompt of only specifying the fruit, the Schur complement component’s KPCA clusters highlight the animals in each cluster (middle case), and similarly by considering the prompt of only specifying the fruit, the animals are captured by each KPCA cluster of the Schur complement matrix. Further decomposition results and analysis are provided in the Appendix.
Moreover, to test how CLIP correction affects the underlying directions of concepts, we applied CLIPDiffusion [30] framework to edit the image according to different CLIP embeddings. Figure 9 illustrates the setup of the problem, where we edit ’initialization image’ that consists of two subjects: a cat and a basketball. Next we denoise and guide the generation according to three different embeddings: Unchanged CLIP embedding of ’initialization image’, corrected CLIP by a ’cat’ direction or corrected CLIP by a ’basketball’ direction. We show that after removing a concept direction, denoiser is no longer rewarded for generating corresponding concept, which is reflected in denoised images. A basketball after correction resembles a bowl with plants and cat loses all of its features. We also note that in both cases, the other object remains intact. To further showcase these results, we performed the same diffusion on the animals with traffic signs dataset shown on the side of Figure 9. This experiment offers a visual interpretation of how the decomposition process removes specific directional components and what the resulting corrected embedding represents. While the $\mathrm{SCE}$-based decomposition is not intended for direct application in image editing tasks, these results are illustrative of the changes that occur in the embedding space when a particular concept is removed from an image. To further validate our approach, we include a baseline where the correction matrix is set to $\Gamma^{*}=\mathrm{I}$, effectively bypassing the optimization step. In this setting, the resulting embeddings fail to preserve the original semantic directions and deviate significantly from the expected outcomes, underscoring the importance of computing the optimal correction matrix $\Gamma^{*}$ for meaningful concept removal. The additional results are provided in the Appendix.
Typographic Attacks. CLIP’s sensitivity to text within images, as showed in [13], makes it vulnerable to typographic attacks, where overlaying misleading text on an image influences its classification [31]. To investigate this, we constructed a dataset of 10 ImageNet classes, each image engraved with a random class label. We conducted an experiment to evaluate the classification performance of CLIP before and after applying our correction method. Figure 10 presents the distribution of predicted classes for images overlaid with the text "cassette player." Prior to correction, CLIP misclassified nearly 79% of these images as "cassette player," indicating a strong susceptibility to typographic bias. However, after applying the $\mathrm{SCE}$-based correction, the misclassification rate drops dramatically to 12%, demonstrating the effectiveness of our approach in mitigating the effect of a typographic attack in the embedding space.
@@
Furthermore, Figure 11 demonstrates CLIP’s susceptibility, showing that top eigenvector directions capture the engraved text rather than the actual image content. By conditioning on prompts like "text reading ’cassette player’ on top of image" and removing this direction, we re-clustered images based on the corrected CLIP embedding. This adjustment allowed us to identify clusters based on the actual image content rather than the engraved text. We visualize the top modes associated with the leading eigenvectors, revealing that the original CLIP embeddings tend to cluster images based on the overlaid text (e.g., text in the center of the image), rather than the actual image content. In contrast, the $\mathrm{SCE}$-corrected CLIP embeddings cluster images according to their visual content, effectively ignoring the spurious text signal.